# Seeing World Dynamics in a Nutshell

### v1

@@ 5.3 Video Downstream Tasks
Video object segmentation. The explicit nature of STAG allows us to propagate object masks in a certain frame to subsequent frames. Specifically, the corresponding STAGs in the object mask can be explicitly selected in the first frame, as demonstrated in previous training free Gaussian segmentation methods [56, 22]. As NutWorld predicts coherent deformation $u_{i}(t)$ for each Gaussian over time, object masks in subsequent frames can be rendered from the STAGs selected initially. In particular, this capability emerges as a by-product without task-specific training [44, 62].
@@ 6 Ablation Study
We analyze NutWorld’s design choices through ablation studies on the 50 selected video clips. As shown in Table 2, our experiments demonstrate that discarding any component from the multi-component pipeline leads to significant performance degradation. Specifically:

### v2

@@ 5.3 Video Downstream Tasks
Video object segmentation. The explicit nature of STAG allows us to propagate object masks in a certain frame to subsequent frames. Specifically, the corresponding STAGs in the object mask can be explicitly selected in the first frame, following previous training free Gaussian segmentation methods [53, 23]. As visualized in following Fig. 6, NutWorld predicts coherent deformation $u_{i}(t)$ for each Gaussian over time, object masks in subsequent frames can be rendered from the STAGs selected initially.
Figure 6: Visualization of Gaussian trajectories. Trajectories of selected Gaussian centers are illustrated as point tracks.
In particular, this capability emerges as a by-product without specific training [59, 70] in video Gaussian representation.
@@ 6 Ablation Study
We analyze NutWorld’s design choices through ablation studies on the 50 selected video clips. As shown in Table 2, our experiments demonstrate that discarding any component from the multi-component pipeline leads to significant performance degradation.